OpenStudy (anonymous):

ALGEBRA 2 help please! What is the simplified form of 4z^2-4z-15/2z^2+z-15? A: 2z+3/z+3 B: 2z-3/z+3 C:2z+3/z-3 D:2z-3/z-3

OpenStudy (campbell_st):

well I'd factorise both quadratics then eliminate common factors so start by factoring the numerator I use this method for a difficult quadratic \[ax^2 + bx + c = 0\] start by multiplying a and c so in your question its 4 x -15 = -60 what are the factors of -60 that add to -4, (hint the larger factor is negative)

OpenStudy (anonymous):

so -10 and 6

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

\[ \frac{4 z^2-4 z-15}{2 z^2+z-15}=\frac{(2 z-5) (2 z+3)}{(z+3) (2 z-5)}=\frac{2 z+3}{z+3} \]